OpenStudy (shamil98):

f(x) = (1/2)^x - 3 How do I graph this? as well as find the domain and range.

OpenStudy (anonymous):

it looks like \(y=\frac{1}{2^x}\) only shifted down 3 units

OpenStudy (anonymous):

domain is all real numbers, because you can raise \(\frac{1}{2}\) to any power you like

OpenStudy (shamil98):

So, the only way to graph this by a graphing calc/wolfram/online calc?

OpenStudy (anonymous):

range is \((-3,\infty)\) because no matter what \(x\) is you know \((\frac{1}{2})^x>0\) and so \[(\frac{1}{2})^x-3>-3\]

OpenStudy (anonymous):

no you don't need wolfram if you have a general idea of what \[y=\frac{1}{2^x}=\left(\frac{1}{2}\right)^x=2^{-x}\] looks like

OpenStudy (anonymous):



OpenStudy (shamil98):

Thank you. Was a bit confused with this. That clarified it.